Operations Management - Learning curve homework

United Research Associates (URA) has received a contract to produce two units of a new cruise missile guidance control. The first unit took 4,000 hours to complete and cost $30,000 in materials and equipment usage. The second took 2,800 hours and cost $18,000 in materials and equipment usage. Labor cost is charged at $22 per hour. The prime contractor has now approached URA and asked to submit a bid for the cost of producing another 20 guidance controls. Use Exhibit 6.4 and Exhibit 6.5. 


a. What will the last unit cost to build? (round your answer to the nearest dollar amount) 

b. What will be the average time for the 20 missile guidance controls? (Round your answer to the nearest whole number)

c. What will the average cost be for guidance control for the 20 in the contract? (Round your answer to the nearest dollar amount) 
Answer #1

ANS a) Final unit cost: material cost + labor cost


Material cost = $18,000


Labor cost = 2800 hours X $22 = 61600


Final unit cost = 18000 + 61600 = $79600


ANS b) 2 unit time = 2800 + 4000 hours = 6800 hours


1 unit time = 6800/2 = 3400


20 unit time = 3400 X 20 = 68000 hours


Average hours = 68000/2 = 34000 hours


ANS c) Total cost of 2 units = material cost of 1 and 2 units + labor cost unit of 1 and 2


Material cost = 30000 + 18000 = 48000


Labor cost = 4000 X 22 + 2800 X 22 = 149600


Total cost of 2 units = 48000 + 149600 = 197600


Total cost of 1 unit = 197600/2 = 98800


Total cost of 20 units = 98800 X 20 = $1976000


answered by: Keynes